# Rethinking Approximate Gaussian Inference in Classification

### v2

@@ C.1 Informal Analysis of Monte Carlo Approximations
We now give an informal theoretical argument for the linear growth of the KL divergence between $\bm{p}$ and $\bm{\hat{P}}^{\lceil S/C\rceil}$ in terms of $C$, under the distributional conditions of the synthetic experiment (Fig. 1).
Taylor expanding Eq. 40 about $\bm{p}$ to second order we obtain
@@
In the synthetic experiment (Fig. 1), the logit class-wise means $\mu_{c}$ and variances $\sigma_{c}^{2}$ are sampled in an i.i.d. way. Let $\bm{Q}\sim\mathcal{N}(\bm{\mu},\operatorname{\mathbf{diag}}(\bm{\sigma}^{2}))$ the unnormalised ‘probabilities’ and $\bm{P}:=\bm{Q}/\sum_{c=1}^{C}Q_{c}$ the probabilities. We have
$\mathrm{Var}[\bm{P}]=\mathbb{E}\left[\frac{\bm{Q}^{2}}{\left(\sum_{c=1}^{C}Q_{c}\right)^{2}}\right]-\mathbb{E}\left[\frac{\bm{Q}}{\sum_{c=1}^{C}Q_{c}}\right]^{2}\approx\frac{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}{\left(\sum_{c=1}^{C}\mathbb{E}[Q_{c}]\right)^{2}}=\frac{\mathrm{Var}[\bm{Q}]}{C^{2}\mathbb{E}[Q_{1}]^{2}}$ (43)
where all operations are taken element-wise. Thus
$\mathbb{E}\left[\|\bm{p}-\bm{P}\|^{2}_{2}\right]=\sum_{c=1}^{C}\mathrm{Var}[P_{c}]\approx\sum_{c=1}^{C}\frac{\mathrm{Var}[Q_{c}]}{C^{2}\mathbb{E}[Q_{1}]^{2}}=\frac{\mathrm{Var}[Q_{1}]}{C\mathbb{E}[Q_{1}]^{2}}.$ (44)
@@
$\mathbb{E}\big{[}\|\bm{p}-\bm{\hat{P}}^{\lceil S/C\rceil}\|^{2}_{2}\big{]}=\frac{1}{\lceil S/C\rceil}\mathbb{E}\left[\|\bm{p}-\bm{P}\|^{2}_{2}\right]\approx\frac{\mathrm{Var}[Q_{1}]}{S\mathbb{E}[Q_{1}]^{2}}.$ (45)
Plugging this into Eq. 42 we get
$\mathbb{E}[\operatorname{D_{KL}}(\bm{p},\bm{\hat{P}}^{\lceil S/C\rceil})]\approx\frac{\mathrm{Var}[Q_{1}]}{2\mathbb{E}[Q_{1}]^{2}}\cdot\frac{C}{S}$ (46)
which grows linearly with the number of classes, as observed in Fig. 1.
C.2 Analysis of the Closed-Form Approximations (Theorem 2.1)
In this section we prove Theorem 2.1 and discuss whether its underlying assumptions are fulfilled in practice. We start with a lemma.
@@
where $\hat{q}$ is the probit approximation (Section A.3). Moreover we write $q_{c}:=q(\mu_{c},\sigma_{c}^{2})$, $\hat{q}_{c}:=\hat{q}(\mu_{c},\sigma_{c}^{2})$, $\bm{p}$ for the true predictive, $\bm{\hat{p}}$ for our approximate predictive, and in addition $\bm{\tilde{p}}$ for the approximate predictive using the exact one dimensional integrals, i.e.
$\bm{p}:=\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}],\quad\bm{\hat{p}}:=\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}},\quad\bm{\tilde{p}}:=\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}.$ (53)
We now restate Theorem 2.1 with an explicit expression for $M(\mathcal{K})$:
@@
as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$
@@
Now for (1), first note that $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$ where [^2] In the case $\varphi=\Phi$ and $Q_{c}\sim\Phi_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$, due to the fast decay of the tail of $\Phi$ we may have $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]=\infty$, making the proof strategy fail in that case. $Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$. By compactness of $\mathcal{K}$, we have in fact $\sup_{(\bm{\mu},\bm{\sigma})\in\mathcal{K}^{C}}\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$. Thus we can apply Lemma C.1 with $(X_{n})_{n\geq 1}$ a sequence of $Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$ and $(Y_{n})_{n\geq 1}$ a sequence of $\sum_{c=1}^{C}Q_{c}$, to get
$|p_{c}-\tilde{p}_{c}|=O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)^{1/2}\right)$ (57)
@@
$(1)$ $=\sum_{c=1}^{C}p_{c}\left(\frac{p_{c}-\tilde{p}_{c}}{p_{c}}+\frac{(p_{c}-\tilde{p}_{c})^{2}}{2\omega(\tilde{p}_{c})^{2}}\right)$ (58)
$=\underbrace{\sum_{c=1}^{C}p_{c}}_{=1}-\underbrace{\sum_{c=1}^{C}\tilde{p}_{c}}_{=1}+\frac{(p_{c}-\tilde{p}_{c})^{2}}{2\omega(\tilde{p}_{c})^{2}}$
$=\frac{O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)^{1/2}\right)^{2}}{2\omega(\tilde{p}_{c})^{2}}$
@@
as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, where $\omega_{c}(\tilde{p}_{c})\in[p_{c},\tilde{p}_{c}]\text{ or }[\tilde{p}_{c},p_{c}]$.
@@
A key assumption in Theorem C.2 is the logit means and variances being restricted to a compact set. In Fig. C.1 we see empirically that this is approximately the case for HET and SNGP models, independently of the number of classes.
Theorem C.2 is of practical value as $M(\mathcal{K}):=\log\left(\frac{1+\delta/u}{1-\Delta}\right)$ is well-behaved:
@@
In other words, given knowledge of the worst case error in the approximation Eq. 15 on the compact set $\mathcal{K}$, we can bound the KL divergence in terms of that error independently of the number of classes. Due to the simplicity of our assumptions, the bound remains quite raw and could be strengthened with further distributional assumptions on the means and variances.
Finally, to obtain a meaningful bound in Theorem C.2, we needed to assume $\Delta(\mathcal{K}):=\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$. In Fig. C.2 we see that this can be assumed to hold on the compact sets on which logit means and variances tend to live (see also Fig. C.1).
Figure C.2: Plot of the ‘true’ predictive $q$ (approximated with a $50000$ sample MC approximation) versus the probit approximation $\hat{q}$, as well as $\frac{q-\hat{q}}{q}$. We see that $\frac{q-\hat{q}}{q}$ tends to be negative, making the assumption $\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$ from Theorem C.2 easily fulfilled.
@@ Appendix D Moment Matching Beta Distributions
As noted in Section 3.2, when $\varphi=\Phi\text{ or }\rho$, we can construct a mapping
@@
where all vector operations are element-wise. Multiplying out the denominators on the right-hand side of the equations of Eq. 60, we obtain a system of two linear equations with two unknowns (for each $c$), which can be solved uniquely, yielding
$\bm{\alpha}$ $:=\frac{\mathbb{E}[\bm{Q}]-\mathbb{E}[\bm{Q}^{2}]}{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}\mathbb{E}[\bm{Q}],$ (61)
$\bm{\beta}$ $:=\frac{\mathbb{E}[\bm{Q}]-\mathbb{E}[\bm{Q}^{2}]}{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}\left(1-\mathbb{E}[\bm{Q}]\right).$
@@ D.1 Information Geometric Interpretation of the Pushforward through NormCDF
Here, we argue that the normCDF activation is an ideal choice for approximating Gaussian pushforwards with Beta distributions by interpreting Fig. D.1. This extends the work from [22], as it shows how one can make sense of the ‘right’ basis for performing Laplace approximations in the classification setting.
@@
The key observation is that $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$ are isometric to $\mathcal{G}(\mathbb{R})$. This is because $\Phi$ and $\rho$ are diffeomorphisms, so in particular sufficient statistics, and the Fisher information metric is invariant under sufficient statistics [2, Section 5.1.4]. Visually, this means that $\varphi_{*}\mathcal{G}(\mathbb{R})$ has the same shape irrespectively of the diffeomorphism activation function $\varphi$. One can thus observe that, given that $\mathcal{B}((0,1))$ is not isometric to $\mathcal{G}(\mathbb{R})$, there exists no activation $\varphi$ such that $\varphi_{*}\mathcal{G}(\mathbb{R})=\mathcal{B}((0,1))$. To design an activation $\varphi$ that maps Gaussians to Betas, the best one can hope to do is to map one specific Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ to a specific Beta distribution $B(\alpha,\beta)$. This is done with the map $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}$ where $\Phi_{\mu,\sigma^{2}}$ and $F_{\alpha,\beta}$ are the cumulative distribution functions of $\mathcal{N}(\mu,\sigma^{2})$ and $B(\alpha,\beta)$ respectively. Taking $\mu=0$, $\sigma^{2}=1$, $\alpha=\beta=1$ we get $\Phi_{0,1}=\Phi$ and $F_{1,1}=\operatorname{id}_{(0,1)}$, yielding $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}=\Phi$.
Now $\mathcal{G}(\mathbb{R})$, and hence $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$, is isometric to the hyperbolic plane Ay et al. [2, Example 3.1]. We can embed this isometrically into Minkowski space with the classical hyperboloid model of the hyperbolic plane [31],
@@
For $\mathcal{B}((0,1))$, we use the isometric embedding from Le Brigant et al. [18, Proposition 2]:
@@
[…]
Finally, we choose our embedding Eq. 62 for $\Phi_{*}\mathcal{G}(\mathbb{R})$ such that it intersects the embedding Eq. 63 at a point, to highlight that the statistical manifolds $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ intersect at a point in the infinite-dimensional ambient space $\mathcal{M}((0,1))$, while $\rho_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ do not.
Moreover, since $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$, our moment matching approximation Eq. 60 is exact when $\mathrm{f}(x)$ is the standard Normal distribution.

### v3

@@ C.1 Informal Analysis of Monte Carlo Approximations
We now give an informal theoretical argument for the linear growth of the KL divergence between $\bm{p}$ and $\bm{\hat{P}}^{\lceil S/C\rceil}$ in terms of $C$, under the distributional conditions of the synthetic experiment (Fig.˜1).
Taylor expanding Eq.˜41 about $\bm{p}$ to second order we obtain
@@
In the synthetic experiment (Fig.˜1), the logit class-wise means $\mu_{c}$ and variances $\sigma_{c}^{2}$ are sampled in an i.i.d. way. Let $\bm{Q}\sim\mathcal{N}(\bm{\mu},\operatorname{\mathbf{diag}}(\bm{\sigma}^{2}))$ the unnormalised ‘probabilities’ and $\bm{P}:=\bm{Q}/\sum_{c=1}^{C}Q_{c}$ the probabilities. We have
$\mathrm{Var}[\bm{P}]=\mathbb{E}\left[\frac{\bm{Q}^{2}}{\left(\sum_{c=1}^{C}Q_{c}\right)^{2}}\right]-\mathbb{E}\left[\frac{\bm{Q}}{\sum_{c=1}^{C}Q_{c}}\right]^{2}\approx\frac{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}{\left(\sum_{c=1}^{C}\mathbb{E}[Q_{c}]\right)^{2}}\approx\frac{\mathrm{Var}[\bm{Q}]}{C^{2}\mathbb{E}[Q_{1}]^{2}}$ (44)
where all operations are taken element-wise, and $\mathbb{E}[Q_{c}]\approx\mathbb{E}[Q_{1}]$ follows from the fact that the $\mu_{c}$ and $\sigma_{c}^{2}$ are i.i.d. Thus
$\mathbb{E}\left[\|\bm{p}-\bm{P}\|^{2}_{2}\right]=\sum_{c=1}^{C}\mathrm{Var}[P_{c}]\approx\sum_{c=1}^{C}\frac{\mathrm{Var}[Q_{c}]}{C^{2}\mathbb{E}[Q_{1}]^{2}}\approx\frac{\mathrm{Var}[Q_{1}]}{C\mathbb{E}[Q_{1}]^{2}}.$ (45)
@@
$\mathbb{E}\big[\|\bm{p}-\bm{\hat{P}}^{\lceil S/C\rceil}\|^{2}_{2}\big]=\frac{1}{\lceil S/C\rceil}\mathbb{E}\left[\|\bm{p}-\bm{P}\|^{2}_{2}\right]\approx\mathbb{E}\left[\frac{\mathrm{Var}[Q_{1}]}{S\mathbb{E}[Q_{1}]^{2}}\right].$ (46)
where the outer expectation is over the (third-order) distributions over $\bm{P}$. Plugging this into Eq.˜43 we get
$\mathbb{E}[\operatorname{D_{KL}}(\bm{p},\bm{\hat{P}}^{\lceil S/C\rceil})]\approx\mathbb{E}\left[\frac{\mathrm{Var}[Q_{1}]}{2\mathbb{E}[Q_{1}]^{2}}\right]\cdot\frac{C}{S}$ (47)
which grows linearly with the number of classes, as observed in Fig.˜1.
C.2 Analysis of the Closed-Form Approximations (Theorem˜3.1)
In this section we prove Theorem˜3.1 and discuss whether its underlying assumptions are fulfilled in practice. We start with a lemma.
@@
where $\hat{q}$ is the probit approximation (Section˜A.3). Moreover we write $q_{c}:=q(\mu_{c},\sigma_{c}^{2})$, $\hat{q}_{c}:=\hat{q}(\mu_{c},\sigma_{c}^{2})$, $\bm{p}$ for the true predictive, $\bm{\hat{p}}$ for our approximate predictive, and in addition $\bm{\tilde{p}}$ for the approximate predictive using the exact one dimensional integrals, i.e.
$\bm{p}:=\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}],\quad\bm{\hat{p}}:=\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}},\quad\bm{\tilde{p}}:=\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}},$ (54)
where $\bm{\Sigma}$ has $\bm{\sigma}$ as diagonal. We now restate Theorem˜3.1 with an explicit expression for $M(\mathcal{K})$:
@@
as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$.
@@
Now for (1), first note that $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$ where [^2] In the case $\varphi=\Phi$ and $Q_{c}\sim\Phi_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$, due to the fast decay of the tail of $\Phi$ we may have $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]=\infty$, and the proof strategy fails in that case. $Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$. By compactness of $\mathcal{K}$, we have in fact $\sup_{(\bm{\mu},\bm{\sigma})\in\mathcal{K}^{C}}\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$. Thus we can apply Lemma˜C.1 with $(X_{n})_{n\geq 1}$ a sequence of $Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$ and $(Y_{n})_{n\geq 1}$ a sequence of $\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}$, to get
$|p_{c}-\tilde{p}_{c}|=O\left(\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{1/2}\right)$ (58)
@@
$(1)$ $=\sum_{c=1}^{C}p_{c}\left(\frac{p_{c}-\tilde{p}_{c}}{p_{c}}+\frac{(p_{c}-\tilde{p}_{c})^{2}}{2\omega_{c}(\tilde{p}_{c})^{2}}\right)$ (59)
$=\underbrace{\sum_{c=1}^{C}p_{c}}_{=1}-\underbrace{\sum_{c=1}^{C}\tilde{p}_{c}}_{=1}+\sum_{c=1}^{C}p_{c}\frac{(p_{c}-\tilde{p}_{c})^{2}}{2\omega_{c}(\tilde{p}_{c})^{2}}$
$=\sum_{c=1}^{C}\frac{p_{c}}{2\omega_{c}(\tilde{p}_{c})^{2}}\cdot O\left(\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{1/2}\right)^{2}$
@@
as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, where $\omega_{c}(\tilde{p}_{c})\in[p_{c},\tilde{p}_{c}]\text{ or }[\tilde{p}_{c},p_{c}]$, where we used that $p_{c}$, $\tilde{p}_{c}$, and hence $\omega_{c}(\tilde{p}_{c})$ is bounded away from 0 by compactness of $\mathcal{K}$.
@@
A key assumption in Theorem˜C.2 is the logit means and variances being restricted to a compact set. In Fig.˜C.1 we see empirically that this is approximately the case for HET and SNGP models, independently of the number of classes.
Theorem˜C.2 is of practical value as $M(\mathcal{K}):=\log\left(\frac{1+\delta/u}{1-\Delta}\right)$ is well-behaved:
@@
In other words, given knowledge of the worst case error in the approximation Eq.˜15 on the compact set $\mathcal{K}$, we can bound the KL divergence in terms of that error independently of the number of classes. Due to the simplicity of our assumptions, the bound remains quite raw and could be strengthened with further distributional assumptions on the means and variances.
Finally, to obtain a meaningful bound in Theorem˜C.2, we needed to assume $\Delta(\mathcal{K}):=\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$. In Fig.˜C.2 we see that this can be assumed to hold on the compact sets on which logit means and variances tend to live (see also Fig.˜C.1).
Figure C.2: Plot of the ‘true’ predictive $q$ (approximated with a $50000$ sample MC approximation) versus the probit approximation $\hat{q}$, as well as $\frac{q-\hat{q}}{q}$. We see that $\frac{q-\hat{q}}{q}$ tends to be negative, making the assumption $\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$ from Theorem˜C.2 easily fulfilled.
Note that, in the proof of Theorem˜C.2, we did not use the form of the probit approximation explicitly, and the theorem holds for any approximation scheme $\hat{q}$ to the one dimensional integrals $q$.
@@ Appendix D Moment Matching Beta Distributions
As noted in Section˜4.2, when $\varphi=\Phi\text{ or }\rho$, we can construct a mapping
@@
where all vector operations are element-wise. Multiplying out the denominators on the right-hand side of the equations of Eq.˜61, we obtain a system of two linear equations with two unknowns (for each $c$), which can be solved uniquely, yielding
$\bm{\alpha}$ $=\frac{\mathbb{E}[\bm{Q}]-\mathbb{E}[\bm{Q}^{2}]}{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}\mathbb{E}[\bm{Q}],$ (62)
$\bm{\beta}$ $=\frac{\mathbb{E}[\bm{Q}]-\mathbb{E}[\bm{Q}^{2}]}{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}\left(1-\mathbb{E}[\bm{Q}]\right).$
@@ D.1 Information Geometric Interpretation of the Pushforward through NormCDF
Here, we argue that the normCDF activation is an ideal choice for approximating Gaussian pushforwards with Beta distributions by interpreting Fig.˜D.1. This extends the work from [mackay_choice_1998], as it shows how one can make sense of the ‘right’ basis for performing Laplace approximations in the classification setting.
@@
The key observation is that $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$ are isometric to $\mathcal{G}(\mathbb{R})$. This is because $\Phi$ and $\rho$ are diffeomorphisms, so in particular sufficient statistics, and the Fisher information metric is invariant under sufficient statistics [ay_information_2017, Section 5.1.4]. Visually, this means that $\varphi_{*}\mathcal{G}(\mathbb{R})$ has the same shape irrespectively of the diffeomorphism activation function $\varphi$. One can thus observe that, given that $\mathcal{B}((0,1))$ is not isometric to $\mathcal{G}(\mathbb{R})$, there exists no activation $\varphi$ such that $\varphi_{*}\mathcal{G}(\mathbb{R})=\mathcal{B}((0,1))$. To design an activation $\varphi$ that maps Gaussians to Betas, the best one can hope to do is to map one specific Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ to a specific Beta distribution $B(\alpha,\beta)$. This is done with the map $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}$ where $\Phi_{\mu,\sigma^{2}}$ and $F_{\alpha,\beta}$ are the cumulative distribution functions of $\mathcal{N}(\mu,\sigma^{2})$ and $B(\alpha,\beta)$ respectively. Taking $\mu=0$, $\sigma^{2}=1$, $\alpha=\beta=1$ we get $\Phi_{0,1}=\Phi$ and $F_{1,1}=\operatorname{id}_{(0,1)}$, yielding $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}=\Phi$.
Now $\mathcal{G}(\mathbb{R})$, and hence $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$, is isometric to the hyperbolic plane ay_information_2017. We can embed this isometrically into Minkowski space with the classical hyperboloid model of the hyperbolic plane [reynolds_hyperbolic_1993],
@@
For $\mathcal{B}((0,1))$, we use the isometric embedding from le_brigant_fisher-rao_2021:
@@
where $\eta(a):=\int_{1}^{a}\sqrt{\psi^{\prime}(r)}\;dr$ and $\psi$ is the digamma function.
Finally, we choose our embedding Eq.˜63 for $\Phi_{*}\mathcal{G}(\mathbb{R})$ such that it intersects the embedding Eq.˜64 at a point, to highlight that the statistical manifolds $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ intersect at a point in the infinite-dimensional ambient space $\mathcal{M}((0,1))$, while $\rho_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ do not.
Moreover, since $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$, our moment matching approximation Eq.˜61 is exact when $\mathrm{f}(x)$ is the standard Normal distribution.